OpenStudy (anonymous):

Use the dartboard below to calculate the expected values in terms of r1=2 bull`s eye worth 3 points r2=4 middle ring worth 1 point r3=2 square root 6 outer ring worth -1 point A) 2/3 points B) 1/6 points C) 16 points D)4/11 points

OpenStudy (anonymous):

Expected value: \[ \Large E[X] = \sum_x xp(x) \]Where \(x\) would be the number of points and \(p( x)\) is the probability of getting \(x\) points.

OpenStudy (anonymous):

what!

OpenStudy (anonymous):

Okay, so you get every possible outcome... you multiply the points you'd get for that outcome by the probability it happens.... then you add all of this up.

OpenStudy (anonymous):

"Use the dartboard below" ??

OpenStudy (anonymous):

Is r1, r2, r3, etc the radius?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay, so what we'll do is calculate the area of each ring. We'll assume that you must hit the dart board and that every point has an equiprobable chance of being hit.

OpenStudy (anonymous):

Then p(x) = Area of x / Total Area

OpenStudy (anonymous):

@Brittanyy! Get it yet?

OpenStudy (anonymous):

kind of im gonna guess it is a

OpenStudy (anonymous):

Don't guess, try

OpenStudy (anonymous):

What is the area of the bulls eye area?

OpenStudy (anonymous):

I would if I understood how you want me to set this problem up to solve it but I have no idea

OpenStudy (anonymous):

Start just by calculating the area of each scoring region. Can you do that?

OpenStudy (anonymous):

so the area for r1 would be 2/3?

OpenStudy (anonymous):

for the bulls eye. I'm getting \(\pi2^2\)

OpenStudy (anonymous):

for the middle ring, I'm getting \(\pi4^2-\pi2^2\)

OpenStudy (anonymous):

and for the outer ring I get \(\pi 2^2 6 - \pi 4^2\) lastly my total area is \(\pi 2^2 6\)

OpenStudy (anonymous):

how are you getting that though?

OpenStudy (anonymous):

Well I'm subtracting the inner circle from the outer circle to get the area of the ring.

OpenStudy (anonymous):

The area of a circle is \(\pi r^2\)

OpenStudy (anonymous):

okay then what do I do?

OpenStudy (anonymous):

So divide the area of each region by the total area of the dart board.